# Spatialyze: A Geospatial Video Analytics System with Spatial-Aware Optimizations

### v3

@@ 6.4.1. High-Level Concepts
Event (ii) is the first frame that the car is no longer visible to the Camera. If this event happens before the car exits its lane, although the car’s trajectory still continues in the lane, the object tracker needs to perform data association at the frame before the car is no longer visible to the Camera as the end of the car’s trajectory.
@@
If a frame contains multiple cars, the Exit Frame Sampler skips to the earliest frame that (i), (ii), or (iii) is estimated to happen for any car, which prevents incorrect tracking for any of them. We implement this idea as our sampling algorithm where given a current frame and its detected cars, it samples the next frame for the object tracker to track. This algorithm starting from the first frame of a video to find the next frame and skips to it. Then it repeats the procedure until the end of the video.
@@ 6.4.2. Sampling Algorithm
Fig. 8 shows the pseudocode of each sampleEvent. exitsLane implements (i) when the car exits the lane. Given the car’s location at the current frame (carLoc) and the lane that contains the car, we assume that the car drives in the lane straight along the lane’s direction. In line LABEL:line:exit-lane-loc, we model the car’s movement as a motion tuple (carLoc, lane.direction). We compute the intersection between the car’s motion tuple and the lane as the location where the car exits the lane. Given the car’s speed v, we know the time it reaches the exit location starting from the current frame in line LABEL:line:exit-lane-time. We sample the frame before the car reaches the exit location. exitsCamera implements (ii) when the car exits the camera’s view. Knowing the car’s current location, moving direction, and speed, carMoves computes its location at any given frame in line LABEL:line:car-moves. In line LABEL:line:view-contain, we use the viewable area of the Camera at each frame derived from § 6.1 to find the first frame when the car’s location is not in the camera’s view. This is the frame when the car already exits the camera’s view, so we sample its preceding frame. newCar implements (iii) when a new car enters the Camera’s view. In line LABEL:line:more-car, we sample the next frame that has more cars detected than the current one. The number of cars is a good heuristic to predict that a new car is entering the frame.
def exitsLane(currentFrame, carLoc, lane, v, cameras):
[…]
def exitsCamera(currentFrame, carLoc, lane, v, cameras):
nextFrame = currentFrame + 1
[…]
cameras[nextFrame].time - cameras[currentFrame].time)): nextFrame += 1
return lastFrameBefore(nextFrame)
def newCar(currentFrame):
nextFrame = currentFrame + 1
[…]
return nextFrame
%**** 3-te.tex Line 100 ****
Figure 8. Computing the events that might trigger a frame sample
At the current frame, the Exit Frame Sampler first finds the frame in which we have more detections for sampleEvent (iii). For each car, the Exit Frame Sampler samples frames for for sampleEvents (i) and (ii). Among (i), (ii), (iii), the Exit Frame Sampler choose the earliest sampleEvent as the next frame for the object tracker to track.
6.4.3. Sampling Algorithm Walk-through
Figure 9. Example Results of our Sampling Algorithm
Figure 10. Sample Events exitsLane (i) and exitsCamera (ii)
@@ 6.4.3. Sampling Algorithm Walk-through
First, we compute the sampleEvent when there are new cars entering $Camera_{B}$’s view through newCar(currentFrame=1). We find that for the rest of the videos, there is only one car in the frame.
We then compute the sampleEvent when $Car_{A}$ exits $Lane_{A}$ with exitsLane(currentFrame=1,carLoc=(70.92,74.7,0), lane=$Lane_{A}$, v=25mph,$cameras$). In line LABEL:line:exit-lane-loc, $Car_{A}$’s motion tuple intersects the polygon of $Lane_{A}$ at location (66.3, 72.7, 0) and $Car_{A}$ is estimated to reach there at 3:51:49.96 PM in line LABEL:line:exit-lane-time. We then find that the last frame before this time is frame 22, depicted as exitsLane in Fig. 10.
Finally, we compute the sampleEvent when the $Car_{A}$ exits the $Camera_{B}$’s view through exitsCamera(currentFrame=1, carLoc=(70.92,74.7,0),lane=$Lane_{A}$,v=25mph,$cameras$). The camera view does not contain $Car_{A}$’s position at frame 35 computed in line LABEL:line:car-moves. Thus, we sample the frame 34, depicted as exitsCamera in Fig. 10.
@@ 6.4.4. Benefits and Limitations
To evaluate the efficacy of the Exit Frame Sampler, we define “skip distance” as the number of frames skipped between two non-skipped frames. We quantify tracking performance using the Exit Frame Sampler by summing the runtime of the sampling algorithm and StrongSORT from the current frame to the next unskipped frame, and divided that by the runtime without running the Exit Frame Sampler. We computed the average for all data points with the same skip distance after processing all the videos from the nuScenes dataset. As Fig. 11 shows, the runtime ratio decreases as the skip distance increases. With an average skip of 3.6 frames, per-frame runtime is reduced to 39% of the original.
Meanwhile, skipping too many frames can result in the tracker losing sight of the target vehicle(s). To assess this risk, we computed the F1 score at each skip distance. Here, “prediction” refers to whether an object that appears in frame $f$ is correctly tracked in frame $(f+\textrm{skip distance})$ as predicted by the Exit Frame Sampler. We compared this with the ground truth where each frame is tracked using StrongSORT without the Exit Frame Sampler. Fig. 11 shows that the F1 score remains robust for skip distances under 14, i.e., accuracy does not decrease much within this threshold. With a skip distance of 13, we observed the runtime to be only 28.27% of the original per frame. Hence, we decided to set the maximum skip distance to be 13 to maximize this accuracy vs runtime tradeoff. However, instances with skip distances over 20 have poor accuracy and they are omitted from the graph to focus on effective metrics.
Figure 11. F1 score and runtime reduction for each skip distance when the Exit Frame Sampler can skip at least 1 frame.
@@ 7. Evaluation
We have implemented a prototype of Spatialyze and evaluated its components against other state-of-the-art systems. We also evaluate our optimization techniques with an ablation study by comparing Spatialyze’s video processing plans with and without each of them. We explain our experiments setup below.
Queries. We implemented Spatialyze’s workflows using four realistic road scenarios, inspired by queries from Scenic (Kim et al., 2021b), VIVA (Romero et al., 2022) and SkyQuery (Bastani et al., 2021). These workflows test Spatialyze against real-life situations. Each workflow starts by loading RoadNetworks and Videos with their Cameras from each query’s corresponding datasets, similar to lines LABEL:line:world-LABEL:line:end-build in Fig. 1. Then, we filter the videos with one of the queries described below, similar to lines LABEL:line:object-LABEL:line:pred-heading. All queries look for objects closer than 50 meters. Finally, we save the filtered videos into video snippets files, similar to line LABEL:line:observe. Table 1 shows the queries in natural language and S-Flow’s filter predicates.
Q1
A pedestrian at an intersection facing perpendicularly to that of a camera. (vr, s)
@@
Q2
2 cars at an intersection moving in opposite directions. (vr, s)
@@
Q3
Camera moving in the direction opposite to the lane direction, with another car (which is moving in the direction of the lane) within 10 meters of it. (vr, s)
contains(lane, [camera, car]) & opposite(lane, camera) & sameDirection(lane, car) & (distance(camera, car) < 10)
Q4
A car and a camera moving in the same direction on lanes; 2 other cars moving together on opposite lanes. (vr, s)
contains(lane1, [car1, camera]) & sameDirection(car1, camera) & contains(lane2, [car2, car3]) & sameDirection(car2, car3) & opposite(lane1, lane2)
Q5
A pedestrian is at an intersection. (vr, s)
contains(intersection, person)
Q6
2 cars are at an intersection. (vr, s)
contains(intersection, [car1, car2])
Q7
A car is on a lane within 10 meters of the camera. (vr, s)
contains(lane, camera) & (distance(camera, car) < 10)
Q8
3 cars, each on a lane. (vr, s)
contains(lane, car1) & contains(lane, car2) & contains(lane, car2)
Q9
A car turning left with a pedestrian at an intersection. (vr, vv)
@@
Q10
A car stopped in a cycling lane. (vr, sq)
contains(bikeLane, car) & stopped(car)
Table 1. Evaluation query descriptions and their S-Flow predicates (VisualRoad [vr]; Scenic [s]; VIVA [vv]; and SkyQuery [sq]).
Dataset. We use the multi-modal AV datasets from Scenic to evaluate Spatialyze. They comprise of 3 datasets that are needed for executing the 4 queries mentioned: 1) videos from on-vehicle cameras, 2) the cameras’ movement and specification, and 3) road network of Boston Seaport where the videos were shot. The former 2 datasets are directly from nuScenes’s training set (Caesar et al., 2019), and the latter is from Scenic (Kim et al., 2021b; Fremont et al., 2023, 2019). From nuScenes, we randomly sampled 80 out of 467 scenes that are captured at the Boston Seaport from on-vehicle cameras; we chose videos from 3 front cameras for each scene; each video is 20 seconds long at 12 Hz. The runtime of Spatialyze’s execution per video does not depend on the number of input video; therefore, we have determined that 240 sampled videos are sufficient to demonstrate our contributions. To compare Spatialyze with VIVA, in addition to the nuScenes dataset, we use the VIVA-provided Jackson Square Traffic Camera dataset which contains 19,469 5-second-long 1080p videos at 30 Hz of a traffic intersection. Finally, to compare with SkyQuery, we use their aerial drone geospatial video datasets, which contains 17,853 frames of a 1080p top-down video, along with the per-frame GPS information.
@@
7.1. Comparison with Competitor Systems
Selecting competitor systems
Since we are not aware of any end-to-end system that focuses on geospatial video analytics, we chose 5 competitor systems to evaluate Spatialyze, where each competitor focuses their novelty of different components covered by Spatialyze.
@@ 7.1. Comparison with Competitor Systems
NuScenes Devkit (Caesar et al., 2019). As the official tool to process NuScenes’ geospatial movement data, we evaluate the runtime-efficiency and the capability of processing large amount of movement objects of the Movable Objects Query Engine against the NuScenes Devkit.
@@
SkyQuery (Bastani et al., 2021). We evaluate Spatialyze against SkyQuery’s geospatial video query processor. As SkyQuery uses entirely different object detector, tracker, and 3d-estimator, we evaluate the video processor’s ability to support different ML functions and how the existing optimization techniques perform with the new ML functions.
7.1.1. EVA (Xu et al., 2022)
We compare the runtime of the Spatialyze pipeline with that of EVA’s implementation using the same queries. EVA is a VDBMS that allows for the definition of custom user-defined functions (UDFs), and also preforms a number of optimizations. Since EVA executes the queries on a frame-by-frame basis, it cannot compute object tracks (and hence object directions), so we use EVA to evaluate only the object detection capability of Spatialyze with Q5-8. To use of EVA’s caching mechanism, we compare with the runtimes of the queries when they are run in series without resetting EVA and its database (so running Q6 directly after Q5, Q7 directly after Q6, etc). The results are shown in Fig. 12. Our goal in comparing against EVA is to evaluate the efficiency of the two system’s query optimization techniques in § 6.1 and § 6.3.
As shown, Spatialyze performs 2-7.3$\times$ faster than EVA on Q5-7, since Spatialyze does not need to run Monodepth2 on each frame, due to the Object Type Pruner and 0pt. In addition, Spatialyze shaves even more time by employing the Road Visibility Pruner to avoid costly YOLO on frames where they are unnecessary. For Q8, our processing time is comparable to EVA. This is because Spatialyze’s Movable Objects Query Engine finds every possible triplets of cars; thus it performs 2 self-joins internally, while EVA only return video frames with 3 cars or more. We also compare the lines of code (LoC) counts between Spatialyze (14-30) and EVA (58-60), showing that Spatialyze makes query implementation easier and requires less LoC for the same results.
@@ 7.1.2. VIVA
Shown in Fig. 12, Spatialyze performs 1.68$\times$ faster compared to VIVA when run on VIVA’s dataset, and 6$\times$ faster when run on the nuScenes dataset. We attribute this increase in performance to Spatialyze’s Object Type Pruner which eliminates the amount of time spent tracking unnecessary objects. VIVA also spends a significantly more time in creating an optimization plan.
In addition, we compare the LoC of the implementations of both VIVA (56) and Spatialyze (51), where Spatialyze provides for a shorter implementation. While we do not have the ground truth to compare with, Spatialyze has higher accuracy in the query results: 45% of tracks returned by Spatialyze are indeed cars turning left. Meanwhile, only 35.7% of the tracks returned by VIVA are correct.
@@
The Devkit implementation of Q4 ran out of memory during execution due to the creation of too many Python objects, as a result of materializing all possible combinations of 3 object tuples in memory before filtering them. Shown in Fig. 12, it also takes significantly longer to execute the same query as Spatialyze, with Spatialyze achieving a 117-716$\times$ faster runtime. Two of the main factors that allow Spatialyze to achieve this better execution time are the following. First, Spatialyze pre-processes the road network data, generating additional columns and indexes that allow it to avoid costly joins, which contribute greatly to the large execution time of Devkit. Second, certain Devkit functions perform costly linear algebra that are avoided by Spatialyze using the geospatial metadata store. Comparing the line counts between the different systems, it took 55-123 lines in Devkit to implement the queries while it only took 12-29 lines to do so in S-Flow.

### v4

@@ 6.4.1. High-Level Concepts
Event (ii) is the first frame that the car becomes invisible to the Camera. If this event happens before the car exits its lane, although its trajectory continues along the lane, the object tracker needs to perform data association at the frame before the car becomes invisible to the Camera as the end of the car’s trajectory.
@@
If a frame contains multiple cars, the Exit Frame Sampler skips to the earliest frame that (i), (ii), or (iii) is estimated to happen for any car, which prevents incorrect tracking for any of them. Our sampling algorithm implements this idea. Given a current frame and its detected cars, it samples the next frame for the object tracker to track. This algorithm starts from the first frame of a video to find the next frame, skips to it, and repeats until the end of the video.
@@ 6.4.2. Sampling Algorithm
§ 6.4.2 shows the pseudocode of each sampleEvent. exitsLane implements (i) when the car exits the lane. Given the car’s location at the current frame (carLoc) and the lane that contains the car, we assume that the car drives in the lane straight along the lane’s direction. In line 6.4.2, we model the car’s movement as a motion tuple (carLoc, lane.direction). We compute the intersection between the car’s motion tuple and the lane as the location where the car exits the lane. Given the car’s speed v, we know the time it reaches the exit location starting from the current frame in line 6.4.2. We sample the frame before the car reaches the exit location. exitsCamera implements (ii) when the car exits the camera’s view. Knowing the car’s current location, moving direction, and speed, carMoves computes its location at any given frame in line 6.4.2. In line 6.4.2, we use the viewable area of the Camera at each frame derived from § 6.1 to find the first frame when the car’s location is not in the camera’s view. This is the frame when the car already exits the camera’s view, so we sample its preceding frame. newCar implements (iii) when a new car enters the Camera’s view. In line 6.4.2, we sample the next frame that has more cars detected than the current one. The number of cars is a good heuristic to predict that a new car is entering the frame.
At the current frame, the Exit Frame Sampler first finds the frame in which we have more detections for sampleEvent (iii). For each car, the Exit Frame Sampler samples frames for sampleEvents (i) and (ii). Among (i), (ii), (iii), the Exit Frame Sampler choose the earliest sampleEvent as the next frame for the object tracker to track.
{listing}
[!ht]{minted}[linenos, mathescape, escapeinside=!!, fontsize=]python def exitsLane(currentFrame, carLoc, lane, v, cameras): exitLaneLoc = intersection((carLoc, lane.direction), lane)!! exitLaneTime = cameras[currentFrame].time + distance(exitLaneLoc, carLoc)/v!! return lastFrameBefore(exitLaneTime) !! def exitsCamera(currentFrame, carLoc, lane, v, cameras): nextFrame = currentFrame + 1 while cameras[nextFrame].view.contains(carMoves(carLoc, lane.direction, v, !! cameras[nextFrame].time - cameras[currentFrame].time)): nextFrame += 1 !! return lastFrameBefore(nextFrame) def newCar(currentFrame): nextFrame = currentFrame + 1 while numCars(nextFrame) ¡= numCars(currentFrame): nextFrame += 1 !! return nextFrame Computing the events that might trigger a frame sample
Sampling Algorithm Walk-through
@@
First, we compute newCar(1) when new cars enter $Camera_{B}$’s view. In the rest of the videos, only one car is in the frame. We then compute exitsLane(1, (70.92,74.7,0), $Lane_{A}$, 25 mph, $cameras$) when $Car_{A}$ exits $Lane_{A}$. In line 6.4.2, $Car_{A}$’s motion tuple intersects the polygon of $Lane_{A}$ at location (66.3, 72.7, 0) at 3:51:49.96 PM in line 6.4.2. The last frame before this time is frame 22, depicted as exitsLane in Fig. 4. Finally, we compute exitsCamera(1, (70.92,74.7,0), $Lane_{A}$, 25 mph, $cameras$) when the $Car_{A}$ exits the $Camera_{B}$’s view. The $Car_{A}$ exits the camera view at frame 35 computed in line 6.4.2. Thus, we sample the frame 34, depicted as exitsCamera in Fig. 4.
@@ 6.4.3. Benefits and Limitations
To evaluate the efficacy of the Exit Frame Sampler, we define “skip distance” as the number of frames skipped between two non-skipped frames. We quantify tracking performance using the Exit Frame Sampler by summing the runtime of the sampling algorithm and StrongSORT from the current frame to the next non-skipped frame and dividing that by the runtime without running the Exit Frame Sampler. After processing all the videos from the nuScenes dataset, we computed the average for all data points with the same skip distance. As Fig. 4 shows, the runtime ratio decreases as the skip distance increases. With an average skip of 3.6 frames, per-frame runtime is 39% of the original.
Meanwhile, skipping too many frames results in the tracker losing sight of the target vehicles. To assess this risk, we computed the F1 score at each skip distance. The ground truth is the object trackings using StrongSORT without the Exit Frame Sampler. Here, “prediction” refers to whether an object at frame $f$ is correctly tracked at frame $(f+\textrm{skip distance})$ as predicted by the Exit Frame Sampler. Fig. 4 shows that the accuracy stays high for skip distances at 13 with the runtime of 28.27% of the original per frame. Hence, we set the maximum skip distance to 13 to maximize this accuracy vs runtime tradeoff. We omit the skip distances over 20 from the graph due to their poor accuracy to focus on effective metrics.
[…]
The figure above is an illustration of exitLane event and exitCamera event. The exitLane event happens earlier. In the figure below, the image shows the first frame that the sampling algorithm starts running on, the frames that the algorithm skips, and the frame that the algorithm skips to (when the car in the first frame exits the lane that it was on). Runtime and Accuracy Evaluation of the Exit Frame Sampler. The left scatter plot shows the object tracking time reduces when the Skip Distance is higher. The right scatter plot shows that the F1 Score (accuracy) of the object tracking algorithm decreases when the Skip Distance is higher. The plot shows our observation that there are more instances of small skipping as a result of the sampling algorithm than instances of large skipping.
[…]
@@ 7. Evaluation
We have implemented a prototype of Spatialyze and evaluated its components against other SOTA systems. We also evaluate our optimization techniques with an ablation study by comparing Spatialyze’s video processing plans with and without each of them. We explain our experiment setup below.
Queries. We implemented Spatialyze’s workflows using four realistic road scenarios, inspired by queries from Scenic (Kim et al., 2021b), VIVA (Romero et al., 2022), and SkyQuery (Bastani et al., 2021). These workflows test Spatialyze against real-life situations. Each workflow starts by loading RoadNetworks and Videos with their Cameras from each query’s corresponding datasets, similar to lines 3-3 in § 3. Then, we filter the videos with one of the queries described in Table 1, similar to lines 3-3. All queries look for objects closer than 50 meters. Finally, we save the filtered videos into video snippet files, similar to line 3. Table 1 shows the queries in natural language and S-Flow’s filter predicates.
Table 1. Evaluation query descriptions and their S-Flow predicates (VisualRoad${}^{\text{\scriptsize[vr]}}$; Scenic${}^{\text{\scriptsize[s]}}$; VIVA${}^{\text{\scriptsize[vv]}}$; and SkyQuery${}^{\text{\scriptsize[vv]}}$).
Q1 A pedestrian at an intersection facing perpendicularly to that of a camera (vr, s).
@@
Q2 2 cars at an intersection moving in opposite directions (vr, s).
@@
Q3 Camera moving in the direction opposite to the lane direction, with another car (which is moving in the direction of the lane) within 10 meters of it (vr, s). contains(lane,[camera, car]) & opposite(lane,camera) & sameDirection(lane,car) & distance(camera,car)<10
Q4 A car and a camera moving in the same direction on lanes; 2 other cars moving together on opposite lanes (vr, s). contains(lane1,[car1,camera]) & sameDirection(car1,camera) & contains(lane2,[car2,car3]) & sameDirection(car2,car3) & opposite(lane1,lane2)
Q5 A pedestrian is at an intersection (vr, s). contains(intersection,person)
Q6 2 cars are at an intersection (vr, s). contains(intersection,[car1,car2])
Q7 A car is on a lane within 10 meters of the camera (vr, s).
contains(lane,camera) & distance(camera,car)<10
Q8 3 cars, each on a lane (vr,s). contains(lane,car1) & contains(lane,car2) & contains(lane,car2)
Q9 A car turning left with a pedestrian at an intersection (vr, vv).
@@
Q10 A car stopped in a cycling lane (vr, sq). contains(bikeLane,car) & stopped(car)
Dataset. We use the multi-modal AV datasets from Scenic (Kim et al., 2021b; Fremont et al., 2023, 2019) to evaluate Spatialyze. They comprise 3 datasets that are needed for executing the 4 queries mentioned: 1) videos from on-vehicle cameras, 2) the cameras’ movement and specification, and 3) the road network of Boston Seaport where the videos were shot. The former 2 datasets are directly from nuScenes’s training set (Caesar et al., 2019), and the latter is from Scenic. From nuScenes, we randomly sampled 80 out of 467 scenes that are captured at the Boston Seaport from on-vehicle cameras; we chose videos from 3 front cameras for each scene; each video is 20 seconds long at 12 FPS. The runtime of Spatialyze’s execution per video does not depend on the number of input videos; therefore, we have determined that 240 sampled videos sufficiently demonstrate our contributions. To compare Spatialyze with VIVA, we use the VIVA-provided Jackson Square Traffic Camera dataset, which contains 19,469 5-second-long 1080p videos at 30 FPS of a traffic intersection, in addition to the nuScenes dataset. Finally, to compare with SkyQuery, we use their aerial drone geospatial video datasets, which contain 17,853 frames of a 1080p top-down video, along with the per-frame GPS information.
@@
7.1. Comparison with Other Systems
Sytems selection
Since we are unaware of any end-to-end system that focuses on geospatial video analytics, we chose 5 systems to evaluate Spatialyze, where we selected each system to evaluate the novelty of different components of Spatialyze.
@@ 7.1. Comparison with Other Systems
NuScenes Devkit (Caesar et al., 2019). As the official tool to process NuScenes’ geospatial movement data, we evaluate the runtime efficiency and the capability of processing large amounts of movement objects of the Movable Objects Query Engine against the NuScenes Devkit.
@@
SkyQuery (Bastani et al., 2021). We evaluate Spatialyze against SkyQuery’s geospatial video query processor. As SkyQuery uses different object detector, tracker, and 3d-estimator, we evaluate the video processor’s ability to support different ML functions and how the existing optimization techniques perform with the new ML functions.
7.1.1. EVA
We compare the runtime of the Spatialyze pipeline with that of EVA’s implementation using the same queries. EVA is a VDBMS that allows custom user-defined functions (UDFs) along with their runtime optimizations. Since EVA executes the queries on a frame-by-frame basis, it cannot compute object tracks (and hence object directions), so we use EVA to evaluate only the object detection capability of Spatialyze with Q5-8. To use EVA’s caching mechanism, we execute the queries on EVA when they are run in series without resetting EVA and its database (running Q6 directly after Q5, Q7 directly after Q6, etc). The results are shown in Fig. 5. Our goal in comparing against EVA is to evaluate the efficiency of the two system’s query optimization techniques in § 6.1 and § 6.3.
As shown, Spatialyze performs 2-7.3$\times$ faster than EVA on Q5-7, since Spatialyze does not need to run Monodepth2 on each frame, due to the Object Type Pruner and 0pt. In addition, Spatialyze shaves even more time by employing the Road Visibility Pruner to avoid costly YOLO on frames where they are unnecessary. For Q8, our processing time is comparable to EVA. This is because Spatialyze’s Movable Objects Query Engine finds every possible triplet of cars; thus it performs 2 self-joins internally, while EVA only returns video frames with 3 cars or more. We also compare the lines of code (LoC) counts between Spatialyze (14-30) and EVA (58-60), showing that Spatialyze makes query implementation easier and requires less LoC for the same results.
@@ 7.1.2. VIVA
Shown in Fig. 5, Spatialyze performs 1.68$\times$ faster compared to VIVA when run on VIVA’s dataset, and 6$\times$ faster when run on the nuScenes dataset. We attribute this increase in performance to Spatialyze’s Object Type Pruner which eliminates the amount of time spent tracking unnecessary objects. VIVA also spends significantly more time creating an optimization plan.
In addition, we compare the LoC of the implementations of VIVA (56) and Spatialyze (51), with Spatialyze providing a shorter implementation. While we do not have the ground truth to compare with, Spatialyze has higher accuracy in the query results: 45% of tracks returned by Spatialyze are indeed cars turning left. Meanwhile, only 35.7% of the tracks returned by VIVA are correct.
@@
The Devkit implementation of Q4 ran out of memory during execution due to the creation of too many Python objects, as a result of materializing all possible combinations of 3 object tuples in memory before filtering them. Shown in Fig. 5, it also takes significantly longer to execute the same query as Spatialyze, with Spatialyze achieving a 117-716$\times$ faster runtime. Two of the main factors that allow Spatialyze to achieve this better execution time are the following. First, Spatialyze pre-processes the road network data, generating additional columns and indexes that allow it to avoid costly joins, which contribute greatly to the large execution time of Devkit. Second, certain Devkit functions perform costly linear algebra that is avoided by Spatialyze using the geospatial metadata store. Comparing the line counts between the different systems, it took 55-123 lines in Devkit to implement the queries while it only took 12-29 lines to do so in S-Flow.